OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin2 x + sin x = 0

OpenStudy (dan815):

hi

OpenStudy (dan815):

is that sin^2 x?

OpenStudy (anonymous):

yes

OpenStudy (dan815):

okk

OpenStudy (anonymous):

sin2 x + sin x = 0 i.e. 2sinx cosx+sinx=0 i.e. sinx (2cosx+1)=0 i.e sinx=0 or (2cosx+1)=0 i.e. x=2npi or cosx = -1/2 = -cos(pi/3)=cos (pi+pi/3)=cos (4pi/3) or x = (n+1) (4pi/3)

OpenStudy (dan815):